OpenStudy (anonymous):

The time between a lightning flash and the following thunderclap may be used to estimate,in kilometers, how far away a storm is. How far away is a storm if 9 seconds elapse between the lightning and the thunder clap? use the formula d=t/3, where t is the time in seconds between the flash and thunderclap. Answer choices:A:27 kilometers B:18 kilometers C:12 kilometers D: 3 kilometers.

OpenStudy (anonymous):

First you take the time between the flash and the clap and divide it by 3 because d=t/3. Plug the 9 into the equation gets us d=9/3.

OpenStudy (anonymous):

okay so that would get me to 3, right? it would be D?

OpenStudy (anonymous):

Yep! 9/3=3 so the storm is 3km away!

OpenStudy (anonymous):

Oh:) haha Okay:D lol thanks so much....

OpenStudy (anonymous):

yay